# Alphazero-like Tree-Search can guide large language model decoding and training

### v1

@@ D.2 SFT and value training details
D.3 Details of value dataset ablation
Here we introduce the details of building mixed, pure and pure,less datasets on GSM8k for value training in Fig 2. For each model checkpoints of 3 epochs during SFT, we first collect 100 outputs per problem in GSM8k training dataset, and then duplicate the overlapped answers, labeled each answer with our training set outcome reward ocracle. we sample multiple output sequences with temperature=0.7, top p=1.0 and top k=100.
@@ D.3 Details of value dataset ablation
For the results in Table 5, the details of training $\{v,\hat{r}\}_{\theta_{1}}$ can be find in Sec D.8. We use MC with $\gamma=1$ to compute the returns. Here we describe the details of training $\{v,\hat{r}\}_{\theta_{1}}^{\text{RL}}$, we use the collected 78.7k samples in Sec D.7 to optimize $\{v,\hat{r}\}_{\theta_{0}}$. The training uses a cosine scheduler decaying from lr=2e-5 to 0.0 with a warmup ratio of 0.03, batch size 128 for 3 epochs.
D.4 Details of applying each tree search approach
@@
$c_{\text{puct}}=\log((\sum_{b}N(s,b)+c_{\text{base}}+1)/c_{\text{base}})+c_{\text{init}}$
In this paper, we fixed $c_{\text{base}}$ with 19652 and set $c_{\text{init}}=3$ for GSM8k, Game24 and RLHF alignment tasks, set $c_{\text{init}}=0.3$ for PrOntoQA tasks. Specifically, in MCTS-$\alpha$, we set the number of simulations before making an action to 5 for GSM8k, Game24 and PrOntoQA, 10 for RLHF alignment. We deterministically sampled actions with the largest visit count. And in MCTS-Rollout, we set an computation upperbound as number of generated tokens or number of model forwards, which is 51200 in GSM8k and Game24, 1800 in PrOntoQA and 5000 in RLHF alignment.
For depth-first search (DFS), the children of a non-leaf node is traversed in a non-decreasing order by value. For efficient exploration, we tried 2 heuristics to prune the subtrees, (1) drop the children nodes with low value by prune_ratio. (2) drop the children nodes lower than a prune_value. We set prune_ratio to be 0.7 for GSM8k, Game24 and PrOntoQA, and 0.9 for RLHF alignment task.
@@
D.5 Details of aggregation results
Here we present the details of how we sample multiple answers with tree search approaches and aggregate them into a final answer.
@@ D.5 Details of aggregation results
In ORM-vote, since we train the ORM with the reward signal -1 and 1, given a list of $N$ answers to be aggregated, we first normalize its values $\{\hat{r}(y^{k})\}_{k=0}^{N-1}$ with min-max normalization make them in $[0,1]$.
D.6 Results per task per aggregation
We show detailed results of GSM8k on Table 7, results of Game24 on Table 8, Table 9 for PrOnToQA and Table 10 for RLHF alignment.
Table 7: Detailed Results in GSM8k
Method N Majority-vote ORM-vote ORM-max #Token
CoT - 41.4 41.4 41.4 0.1k
CoT-SC 1 38.21 38.21 38.21 0.1k
CoT-SC 10 51.93 57.47 53.83 1k
CoT-SC 20 54.44 59.44 54.74 2k
CoT-SC 50 56.79 61.03 54.44 5k
CoT-SC 100 58.15 62.70 53.68 10k
CoT-SC-Tree 1 37.91 37.91 37.91 0.1k
CoT-SC-Tree 10 50.19 53.15 50.95 0.8k
CoT-SC-Tree 20 52.69 55.12 52.84 1.3k
CoT-SC-Tree 50 54.51 57.16 53.45 2.7k
BFS 1 52.5 52.5 52.5 0.5k
BFS 10 58.98 56.25 54.97 3.1k
BFS 20 58.91 56.79 52.39 5.3k
BFS 50 59.29 59.36 53.22 10.1k
DFS 1 51.8 51.8 51.8 0.5k
DFS 10 53.75 54.66 53.37 1.2k
DFS 20 54.89 55.50 53.53 1.6k
MCTS-$\alpha$ (no sampling) 1 51.9 51.9 51.9 0.5k
MCTS-$\alpha$-intra tree 1 46.78 46.78 46.78 0.7k
MCTS-$\alpha$-intra tree 10 57.85 56.86 54.36 3.4k
MCTS-$\alpha$-intra tree 20 58.83 58.23 55.19 5.3k
MCTS-$\alpha$-inter trees 1 51.9 51.9 51.9 0.5k
MCTS-$\alpha$-inter trees 10 57.92 58.53 55.34 5.5k
MCTS-$\alpha$-inter trees 20 58.83 59.06 54.97 11.1k
MCTS-$\alpha$-inter trees 50 58.76 61.26 53.98 27.8k
MCTS 1 52.2 52.2 52.2 0.5k
MCTS 10 57.92 55.72 53.75 2.4k
MCTS 20 58.61 56.79 54.74 4.0k
MCTS 50 59.36 58.23 53.75 7.5k
MCTS-Rollout 1 47.8 47.8 47.8 3.4k
MCTS-Rollout 10 51.10 50.49 49.81, 5.4k
MCTS-Rollout 20 51.86 51.25 50.19 6.1k
MCTS-Rollout 50 52.69 52.24 50.49 7.2k
Table 8: Detailed Results in Game24
Method N Majority-vote ORM-vote ORM-max #Token
CoT - 12.7 12.7 12.7 0.1k
CoT-SC 1 9.94 9.94 9.94 0.1k
CoT-SC 10 13.54 50.83 50.83 0.8k
CoT-SC 20 14.36 65.75 65.47 1.6k
CoT-SC 50 16.30 78.45 78.45 4.0k
CoT-SC 100 18.23 84.25 84.53 7.9k
CoT-SC-Tree 1 9.67 9.67 9.67 0.1k
CoT-SC-Tree 10 11.33 48.34 48.34 0.7k
CoT-SC-Tree 20 13.26 61.60 62.15 1.1k
CoT-SC-Tree 50 16.57 69.61 69.89 2.0k
BFS 1 64.8 64.8 64.8 0.4k
BFS 10 47.79 70.72 70.99 1.6k
BFS 20 27.62 69.34 69.34 2.3k
BFS 50 7.18 70.17 70.72 3.7k
DFS 1 66.3 66.3 66.3 0.4k
DFS 10 50.83 62.98 64.09 0.4k
DFS 20 50.83 62.98 64.09 0.6k
MCTS-$\alpha$ (no sampling) 1 63.3 63.3 63.3 0.4k
MCTS-$\alpha$-intra tree 1 64.36 64.36 64.36 0.4k
MCTS-$\alpha$-intra tree 10 66.85 67.68 63.90 0.9k
MCTS-$\alpha$-intra tree 20 67.13 69.34 68.78 1.1k
MCTS-$\alpha$-intra tree 50 67.96 69.89 69.34 1.4k
MCTS-$\alpha$-inter trees 1 63.3 63.3 63.3 0.4k
MCTS-$\alpha$-inter trees 10 72.65 82.87 82.32 4.1k
MCTS-$\alpha$-inter trees 20 72.93 84.25 83.15 8.3k
MCTS 1 64.0 64.0 64.0 0.4k
MCTS 10 70.44 70.72 70.17 0.8k
MCTS 20 72.10 72.10 71.27 1.1k
MCTS 50 72.38 72.38 71.55 1.6k
MCTS-Rollout 1 71.3 71.3 71.3 0.7k
MCTS-Rollout 10 73.48 73.20 72.65 0.9k
MCTS-Rollout 20 73.48 73.48 72.38 1.0k
MCTS-Rollout 50 73.48 73.48 72.38 1.1k
Table 9: Detailed Results in PrOntoQA
Method N Majority-vote ORM-vote ORM-max #Token
CoT - 48.8 48.8 48.8 77
CoT-SC 1 54.40 54.40 54.40 91.25
CoT-SC 3 63.60 82.40 82.40 273.75
CoT-SC 10 58.40 97.80 97.80 912.55
CoT-SC 20 57.00 99.80 99.80 1.8k
CoT-SC-Tree 1 50.20 50.20 50.20 82.02
CoT-SC-Tree 10 62.40 98.40 98.40 413.58
CoT-SC-Tree 20 61.00 99.40 99.40 632.91
BFS 1 94.40 94.40 94.40 125.52
BFS 10 99.00 100.00 100.00 837.78
BFS 20 98.60 99.80 99.80 1.5k
DFS 1 93.30 93.30 93.30 124.46
DFS 10 95.60 96.20 96.00 178.79
DFS 20 95.60 96.20 96.20 180.92
MCTS-$\alpha$ (no sampling) 1 99.40 99.40 99.40 183.66
MCTS-$\alpha$-intra tree 1 97.20 97.20 97.20 208.68
MCTS-$\alpha$-intra tree 10 99.80 99.80 99.80 364.96
MCTS-$\alpha$-intra tree 20 99.80 99.80 99.80 441.31
MCTS-$\alpha$-inter trees 1 99.40 99.40 99.40 183.66
MCTS-$\alpha$-inter trees 10 100.00 100.00 100.00 1.9k
MCTS-$\alpha$-inter trees 20 100.00 100.00 100.00 3.8k
MCTS 1 94.20 94.20 94.20 126.65
MCTS 10 99.60 99.60 99.60 182.88
MCTS 20 100.00 100.00 100.00 240.16
MCTS-Rollout 1 96.90 96.90 96.90 210.41
MCTS-Rollout 10 99.20 99.20 99.20 220.16
MCTS-Rollout 20 99.20 99.20 99.20 224.16
Table 10: Detailed Results in RLHF alignment
Method N Mean Best #Forward
CoT 1 0.387 0.387 57.8
CoT-SC 1 -0.164 -0.164 58
CoT-SC 10 -0.182 1.592 0.6k
CoT-SC 20 -0.175 1.972 1.2k
CoT-SC 50 -0.176 2.411 2.9k
BFS 1 -1.295 -1.295 61.8
BFS 10 -1.523 -1.065 0.6k
BFS 20 -1.520 -0.948 1.2k
BFS 50 -1.474 -0.813 3.1k
DFS 1 -1.295 -1.295 61.8
DFS 10 -1.67 -1.17 64.8
DFS 20 -1.69 -1.08 66.9
DFS 50 -1.71 -0.99 71.9
MCTS-$\alpha$ (no sampling) 1 2.221 2.221 186
MCTS-$\alpha$-intra tree 1 1.538 1.538 198.50
MCTS-$\alpha$-intra tree 10 1.527 3.052 1.6k
MCTS-$\alpha$-intra tree 20 1.533 3.311 3.1k
MCTS 1 -1.295 -1.295 61.8
MCTS 10 -1.146 0.160 0.6k
MCTS 20 -1.08 0.528 1.2k
MCTS 50 -0.961 0.981 2.8k
MCTS-Rollout 1 1.925 1.925 0.8k
MCTS-Rollout 10 2.278 2.540 1.1k
MCTS-Rollout 20 2.376 2.643 1.2k
MCTS-Rollout 50 2.491 2.746 1.3k
D.7 sampling details of iterative update
@@ D.6 Results per task per aggregation
D.8 Training details of iterative update
Policy training in GSM8k: We construct the dataset for supervised finetuning by combining data in the training dataset with 57.6k correct answers sampled in Sec D.7 which results in 64.1k distinct correct answers. And we train the new policy model $\pi_{\theta_{1}}$ from the starting base model LLaMA2-7b for 3 epochs, following Yuan et al. (2023a). The training setting is the same as described in Sec D.2.
Value training in GSM8k: We construct the dataset for value and ORM training by combining the data used to train $\{v,\hat{r}\}_{\theta_{0}}$ with 78.7k answers sampled by MCTS-$\alpha$ in Sec D.7. To fairly compare $\{v,\hat{r}\}_{\theta_{1}}$ with $\{v,\hat{r}\}_{\theta_{0}}$, we drop samples in the former dataset to keep at most $51-12=39$ answers per problem resulting in 359k distinct answers. And we train the new value function $\{v,\hat{r}\}_{\theta_{1}}$ from the value model with its initial weight(before being updated on any data) for 3 epochs. The training setting is the same as described in Sec D.2.
Policy training in RLHF alignment: For the MCTS-$\alpha$’s training, we subsample the top 5 answers from the full 10 candidates (mentioned in Appendix D.7) to serve as the SFT dataset. For the RFT n=5 baseline, we subsample the top 5 answers from 50 direct decodings as the SFT dataset. For the training of the PPO algorithm, we adopt the implementation from trlx [^4] https://github.com/CarperAI/trlx. We sample 20 answers for each question in total, which maintains the same level of token consumption during the PPO rollouts as that of MCTS-$\alpha$.
Value training in RLHF alignment: We construct the value and ORM dataset by mixing data from SFT-policy direct decoding and from MCTS-$\alpha$. To make the comparison fair, the new value function’s training utilizes the same amount of data as the old one by subsampling 40 answers (from 50 shown in Appendix D.2) from direct decoding data and all 10 answers (shown in Appendix D.7) generated by MCTS-$\alpha$. We train our value function with learning rate 2e-5 and cosine scheduler from the initial model (instead of continuing training from the old value function) for 2 epochs.
Appendix E Qualitative Results
Here we present some qualitative results of generation for different methods (Path@1 performance) in Table 13 (for Game24) and 14 (for the RLHF alignment task).
Prompt: 4 8 9 13

### v2

@@ E.2 SFT and value training details
SFT in Chess Endgame: Currently we use the opensourced behavior-cloning model in Abdulhai et al. (2023) as the SFT model [^4] https://github.com/abdulhaim/LMRL-Gym. Given a Forsyth-Edwards Notation (FEN) state description, the language model policy autoregressively outputs SAN(Standard Algebraic Notation)-format actions.
Value training in Chess Endgame Currently we use the opensourced critic value function model trained by Proximal Policy Optimization (PPO) in Abdulhai et al. (2023). The language model value function receives a FEN state description and outputs a scalar as the value estimation.
E.3 Details of comparing prompting-based value function and learned value function
For the prompting-based value function presented in Table 2, we use GPT-3.5-0613 for GPT3.5 model and design few-shot prompts for GPT3.5-policy/value and also LLaMA2-7B. We sample 1 time from GPT3.5 and 3 times from LLaMA2-7B to form the value evaluation.
E.4 Details of value dataset ablation
Here we introduce the details of building mixed, pure and pure,less datasets on GSM8k for value training in Table 5. For each model checkpoints of 3 epochs during SFT, we first collect 100 outputs per problem in GSM8k training dataset, and then duplicate the overlapped answers, labeled each answer with our training set outcome reward ocracle. we sample multiple output sequences with temperature=0.7, top p=1.0 and top k=100.
@@ E.4 Details of value dataset ablation
For the results in Table 7, the details of training $\{v,\hat{r}\}_{\phi_{1}}$ can be find in Sec E.9. We use MC with $\gamma=1$ to compute the returns. Here we describe the details of training $\{v,\hat{r}\}_{\phi_{1}}^{\text{RL}}$, we use the collected 78.7k samples in Sec E.8 to optimize $\{v,\hat{r}\}_{\phi_{0}}$. The training uses a cosine scheduler decaying from lr=2e-5 to 0.0 with a warmup ratio of 0.03, batch size 128 for 3 epochs.
E.5 Hyperparameter Selection Protocols
Here we present the selection protocols of hyperparameters.
[…]
Specific hyperparameters for Monte Carlo Tree Search variants: Basically we adopted the default values from Schrittwieser et al. (2020b) and Silver et al. (2017) for most of the hyperparameters, such as $c_{\text{base}}=19652$ in Equation 4, $\tau=1.0$ for MCTS-alpha stochastic search. And for the Dirichlet noise of MCTS-$\alpha$ stochastic search as mentioned in Appendix E.7, we adopted the default value in Silver et al. (2017) as 0.3, which is specified for chess. We do find that in MCTS-$\alpha$, MCTS-Rollout and MCTS, $c_{\text{init}}$ can affect the balance between exploration and exploitation, and we chose it by running several trials from two possible values: $\{0.3,3.0\}$. Moreover, for MCTS-$\alpha$, the hyperparameter num of simulation, $n_{\text{simulation}}$ is chosen as 5 for shallow trees (tree max-depths less than of equal to 15 over GSM8k, Game24 and ProntoQA) and 10 in deep trees(a tree max-depth of 64 in RLHF), controlling the search complexity at each step.
Specific hyperparameters for BFS-/DFS-V: BFS-V does not have hyperparameters for single search. For DFS-V, the children of a non-leaf node is traversed in a non-decreasing order by its value. For the sake of efficient exploration, we tried 2 heuristics to prune the subtrees, (1) drop the children nodes with lower values by prune_ratio. (2) drop the children nodes lower than a prune_value. The latter is adopted from Yao et al. (2023). In our experiments, we tried possible prune_values from $\{0.5,0.0,-0.5\}$ or None, we found that setting a high prune_value like $0.5$ or $0.0$ may introduce significant performance drop, however, setting a higher prune_value may introduce very closer answers. Therefore, we finally use prune_ratio for efficient exploration during searching on the tree with DFS-V. We set prune_ratio to be $0.7$ (selected from $\{0.3,0.5,0.7\}$) for GSM8k, Game24 and PrOntoQA (tree-max-widths of 6, 6, 20), and 0.95 (selected from $\{0.5,0.7,0.9,0.9\}$)for RLHF alignment task since its much wider(a tree-max-width of 50).
E.6 Details of applying each tree search approach
@@
$c_{\text{puct}}=\log((\sum_{b}N(s,b)+c_{\text{base}}+1)/c_{\text{base}})+c_{\text{init}}$ (4)
In this paper, we fixed $c_{\text{base}}$ with 19652 and set $c_{\text{init}}=3$ for GSM8k, Game24 and RLHF alignment tasks, set $c_{\text{init}}=0.3$ for PrOntoQA tasks. For Chess Endgame, we manually set $c_{\text{puct}}=0$. Specifically, in MCTS-$\alpha$, we set the number of simulations before making an action to 5 for GSM8k, Game24 and PrOntoQA, 10 for RLHF alignment and Chess Endgame. We deterministically sampled actions with the largest visit count. And in MCTS-Rollout, we set an computation upperbound as number of generated tokens or number of model forwards, which is 51200 in GSM8k and Game24, 1800 in PrOntoQA and 5000 in RLHF alignment, 10000 in Chess Endgame.
For DFS-V, the children of a non-leaf node is traversed in a non-decreasing order by value. For efficient exploration, we tried 2 heuristics to prune the subtrees, (1) drop the children nodes with low value by prune_ratio. (2) drop the children nodes lower than a prune_value. We set prune_ratio to be 0.7 for GSM8k, Game24 and PrOntoQA, and 0.95 for RLHF alignment task, 0.6 for Chess Endgame.
@@
E.7 Details of aggregation experiments
Another alternative setting for conducting multiple searches in Inter-tree Search. Inter-tree Search builds a new tree for each new search, which increases the diversity of the search space, with extra computation burdens proportional to the search times. Thus, the intra-setting will have a larger state space compared with intra-tree setting. Our experiment results shown in Fig 2 (comparing MCTS-$\alpha$ intra-tree and inter-tree settings) also verify the performance gain brought by the larger search space.
We also present the details of how we sample multiple answers with tree search approaches and aggregate them into a final answer.
For the results of CoT-SC-Tree on Table 3 and Table 6, they can be viewed as intra-tree searches. For the results in Figure 2, only MCTS-$\alpha$ inter-trees were conducted with inter-tree searches, other tree-search algorithms (MCTS, MCTS-Rollout, BFS-V, DFS-V) were all conducted with intra-tree searches
Note that for all tree search algorithms except BFS-V, multiple searches are conducted in a sequential manner, while for BFS-V which can actually be regarded as Beam-Search, the number of searches means the number of beam size.
@@ E.7 Details of aggregation experiments
In ORM-vote, since we train the ORM with the reward signal -1 and 1, given a list of $N$ answers to be aggregated, we first normalize its values $\{\hat{r}(y^{j})\}_{j}$ with min-max normalization make them in $[0,1]$.
E.8 sampling details of iterative update
@@
E.9 Training details of iterative update
Policy training in GSM8k: We construct the dataset for supervised finetuning by combining data in the training dataset with 57.6k correct answers sampled in Sec E.8 which results in 64.1k distinct correct answers. And we train the new policy model $\pi_{\theta_{1}}$ from the starting base model LLaMA2-7b for 3 epochs, following Yuan et al. (2023a). The training setting is the same as described in Sec E.2.
Value training in GSM8k: We construct the dataset for value and ORM training by combining the data used to train $\{v,\hat{r}\}_{\phi_{0}}$ with 78.7k answers sampled by MCTS-$\alpha$ in Sec E.8. To fairly compare $\{v,\hat{r}\}_{\phi_{1}}$ with $\{v,\hat{r}\}_{\phi_{0}}$, we drop samples in the former dataset to keep at most $51-12=39$ answers per problem resulting in 359k distinct answers. And we train the new value function $\{v,\hat{r}\}_{\phi_{1}}$ from the value model with its initial weight(before being updated on any data) for 3 epochs. The training setting is the same as described in Sec E.2.
Policy training in RLHF alignment: For the MCTS-$\alpha$’s training, we subsample the top 5 answers from the full 10 candidates (mentioned in Appendix E.8) to serve as the SFT dataset. For the RFT n=5 baseline, we subsample the top 5 answers from 50 direct decodings as the SFT dataset. For the training of the PPO algorithm, we adopt the implementation from trlx [^5] https://github.com/CarperAI/trlx. We sample 20 answers for each question in total, which maintains the same level of token consumption during the PPO rollouts as that of MCTS-$\alpha$.
Value training in RLHF alignment: We construct the value and ORM dataset by mixing data from SFT-policy direct decoding and from MCTS-$\alpha$. To make the comparison fair, the new value function’s training utilizes the same amount of data as the old one by subsampling 40 answers (from 50 shown in Appendix E.2) from direct decoding data and all 10 answers (shown in Appendix E.8) generated by MCTS-$\alpha$. We train our value function with learning rate 2e-5 and cosine scheduler from the initial model (instead of continuing training from the old value function) for 2 epochs.
Appendix F Qualitative Results
Here we present some qualitative results of generation for different methods in Table 18 (for Game24), Table 19 (for GSM8k), Table 20 (for ProntoQA) and 21 (for the RLHF alignment task).
Problem: 4 8 9 13
@@ Methods
Question: A new program had 60 downloads in the first month. The number of downloads in the second month was three times as many as the downloads in the first month, but then reduced by 30% in the third month. How many downloads did the program have total over the three months?
Methods
Response
Correct
CoT
The number of downloads in the second month was 3*60 = 180
The number of downloads in the third month was 180*0.30 = 54
The total number of downloads for the three months was 60+180+54 = 294
The answer is 294
No
BFS-V
In the second month, the program had 60*3=180 downloads
In the third month, the program had 180*30/100=54 downloads which was 30% less than the previous month
Total downloads for the program in the three months was 60+180+54=294
The answer is 294
No
MCTS-$\alpha$
The program had 60 x 3 = 180 downloads in the second month.
The downloads in the third month was 180 x 0.30 = 54 downloads lesser than the previous month.
So, the program had 180 - 54 = 126 downloads in the third month.
Therefore, the program had 60 + 180 + 126 = 366 downloads over the three months.
The answer is 366
Yes
MCTS-Rollout
The number of downloads in the second month was 3 times as many as the downloads in the first month, which means the program had 3*60 = 180 downloads in the second month.
The total number of downloads for the two months was 180+60 = 240 In the third month, the program had 180-30/100*180 = 120 downloads, 30% of the number of downloads it had in the second month.
The total number of downloads for the three months was 240+120 = 360
The answer is 360
No
Table 19: Qualitative examples of GSM8k outputs
Question: Butterflies are lepidopterans. Every arthropod is small. Whales are not small. Invertebrates are animals. Every insect is an arthropod. Lepidopterans are insects. Every insect is six-legged. Every arthropod is an invertebrate. Animals are multicellular. Polly is a lepidopteran. Is the statement ”Polly is not small” true or false?
Methods
Response
Correct
CoT
Polly is a lepidopteran.
Lepidopterans are insects.
Polly is an insect.
Every insect is an arthropod.
Polly is an arthropod.
Every arthropod is small.
Polly is small.
The answer is true.
No
BFS-V
Polly is a lepidopteran.
Lepidopterans are insects.
Polly is an insect.
Every insect is an arthropod.
Polly is an arthropod.
Every arthropod is small.
Polly is small.
Polly is small.
The answer is false.
Yes
MCTS-$\alpha$
Polly is a lepidopteran.
Lepidopterans are insects.
Polly is an insect.
Every insect is an arthropod.
Polly is an arthropod.
Every arthropod is small.
Polly is small.
The answer is false.
Yes
MCTS-Rollout
Polly is a lepidopteran.
Lepidopterans are insects.
Polly is an insect.
Every insect is an arthropod.
Polly is an arthropod.
Every arthropod is small.
Polly is small.
The answer is false.
Yes
Table 20: Qualitative examples of ProntoQA outputs
@@
Appendix G Detailed Experiment Results
G.1 Results per task per aggregation
We show detailed results of GSM8k on Table 22, results of Game24 on Table 23, Table 24 for PrOnToQA and Table 25 for RLHF alignment. Due to the limit of computation resources, we show the results under 1 seed except for the path@1 results.
Table 22: Detailed Results in GSM8k
Method N Majority-vote ORM-vote ORM-max #Token
CoT - 41.4 41.4 41.4 0.1k
CoT-SC 1 38.21 38.21 38.21 0.1k
CoT-SC 10 51.93 57.47 53.83 1k
CoT-SC 20 54.44 59.44 54.74 2k
CoT-SC 50 56.79 61.03 54.44 5k
CoT-SC 100 58.15 62.70 53.68 10k
CoT-SC-Tree 1 37.91 37.91 37.91 0.1k
CoT-SC-Tree 10 50.19 53.15 50.95 0.8k
CoT-SC-Tree 20 52.69 55.12 52.84 1.3k
CoT-SC-Tree 50 54.51 57.16 53.45 2.7k
BFS-V 1 52.5 52.5 52.5 0.5k
BFS-V 10 58.98 56.25 54.97 3.1k
BFS-V 20 58.91 56.79 52.39 5.3k
BFS-V 50 59.29 59.36 53.22 10.1k
DFS-V 1 51.8 51.8 51.8 0.5k
DFS-V 10 57.09 56.18 54.89 1.2k
DFS-V 20 58.23 58.38 55.19 1.6k
DFS-V 50 58.98 58.98 55.35 2.1k
MCTS-$\alpha$ (no sampling) 1 51.9 51.9 51.9 0.5k
MCTS-$\alpha$-intra tree 1 46.78 46.78 46.78 0.7k
MCTS-$\alpha$-intra tree 10 57.85 56.86 54.36 3.4k
MCTS-$\alpha$-intra tree 20 58.83 58.23 55.19 5.3k
MCTS-$\alpha$-inter trees 1 51.9 51.9 51.9 0.5k
MCTS-$\alpha$-inter trees 10 57.92 58.53 55.34 5.5k
MCTS-$\alpha$-inter trees 20 58.83 59.06 54.97 11.1k
MCTS-$\alpha$-inter trees 50 58.76 61.26 53.98 27.8k
MCTS 1 52.2 52.2 52.2 0.5k
MCTS 10 57.92 55.72 53.75 2.4k
MCTS 20 58.61 56.79 54.74 4.0k
MCTS 50 59.36 58.23 53.75 7.5k
MCTS-Rollout 1 47.8 47.8 47.8 3.4k
MCTS-Rollout 10 51.10 50.49 49.81, 5.4k
MCTS-Rollout 20 51.86 51.25 50.19 6.1k
MCTS-Rollout 50 52.69 52.24 50.49 7.2k
Table 23: Detailed Results in Game24
Method N Majority-vote ORM-vote ORM-max #Token
CoT - 12.7 12.7 12.7 0.1k
CoT-SC 1 9.94 9.94 9.94 0.1k
CoT-SC 10 13.54 50.83 50.83 0.8k
CoT-SC 20 14.36 65.75 65.47 1.6k
CoT-SC 50 16.30 78.45 78.45 4.0k
CoT-SC 100 18.23 84.25 84.53 7.9k
CoT-SC-Tree 1 9.67 9.67 9.67 0.1k
CoT-SC-Tree 10 11.33 48.34 48.34 0.7k
CoT-SC-Tree 20 13.26 61.60 62.15 1.1k
CoT-SC-Tree 50 16.57 69.61 69.89 2.0k
BFS-V 1 64.8 64.8 64.8 0.4k
BFS-V 10 47.79 70.72 70.99 1.6k
BFS-V 20 27.62 69.34 69.34 2.3k
BFS-V 50 7.18 70.17 70.72 3.7k
DFS-V 1 66.3 66.3 66.3 0.4k
DFS-V 10 55.25 69.06 69.34 0.9k
DFS-V 20 54.14 69.34 69.61 1.0k
MCTS-$\alpha$ (no sampling) 1 63.3 63.3 63.3 0.4k
MCTS-$\alpha$-intra tree 1 64.36 64.36 64.36 0.4k
MCTS-$\alpha$-intra tree 10 66.85 67.68 63.90 0.9k
MCTS-$\alpha$-intra tree 20 67.13 69.34 68.78 1.1k
MCTS-$\alpha$-intra tree 50 67.96 69.89 69.34 1.4k
MCTS-$\alpha$-inter trees 1 63.3 63.3 63.3 0.4k
MCTS-$\alpha$-inter trees 10 72.65 82.87 82.32 4.1k
MCTS-$\alpha$-inter trees 20 72.93 84.25 83.15 8.3k
MCTS 1 64.0 64.0 64.0 0.4k
MCTS 10 70.44 70.72 70.17 0.8k
MCTS 20 72.10 72.10 71.27 1.1k
MCTS 50 72.38 72.38 71.55 1.6k
MCTS-Rollout 1 71.3 71.3 71.3 0.7k
MCTS-Rollout 10 73.48 73.20 72.65 0.9k
MCTS-Rollout 20 73.48 73.48 72.38 1.0k
MCTS-Rollout 50 73.48 73.48 72.38 1.1k
Table 24: Detailed Results in PrOntoQA
Method N Majority-vote ORM-vote ORM-max #Token
CoT - 48.8 48.8 48.8 92
CoT-SC 1 54.40 54.40 54.40 91.25
CoT-SC 3 63.60 82.40 82.40 273.75
CoT-SC 10 58.40 97.80 97.80 912.55
CoT-SC 20 57.00 99.80 99.80 1.8k
CoT-SC-Tree 1 50.20 50.20 50.20 82.02
CoT-SC-Tree 10 62.40 98.40 98.40 413.58
CoT-SC-Tree 20 61.00 99.40 99.40 632.91
BFS-V 1 94.40 94.40 94.40 125.52
BFS-V 10 99.00 100.00 100.00 837.78
BFS-V 20 98.60 99.80 99.80 1.5k
DFS-V 1 93.30 93.30 93.30 124.46
DFS-V 10 95.60 96.40 96.40 187.59
DFS-V 20 95.60 96.40 96.40 193.91
MCTS-$\alpha$ (no sampling) 1 99.40 99.40 99.40 183.66
MCTS-$\alpha$-intra tree 1 97.20 97.20 97.20 208.68
MCTS-$\alpha$-intra tree 10 99.80 99.80 99.80 364.96
MCTS-$\alpha$-intra tree 20 99.80 99.80 99.80 441.31
MCTS-$\alpha$-inter trees 1 99.40 99.40 99.40 183.66
MCTS-$\alpha$-inter trees 10 100.00 100.00 100.00 1.9k
MCTS-$\alpha$-inter trees 20 100.00 100.00 100.00 3.8k
MCTS 1 94.20 94.20 94.20 126.65
MCTS 10 99.60 99.60 99.60 182.88
MCTS 20 100.00 100.00 100.00 240.16
MCTS-Rollout 1 96.90 96.90 96.90 210.41
MCTS-Rollout 10 99.20 99.20 99.20 220.16
MCTS-Rollout 20 99.20 99.20 99.20 224.16
Table 25: Detailed Results in RLHF alignment
Method N Mean Best #Forward
CoT 1 0.387 0.387 57.8
CoT-SC 1 -0.164 -0.164 58
CoT-SC 10 -0.182 1.592 0.6k
CoT-SC 20 -0.175 1.972 1.2k
CoT-SC 50 -0.176 2.411 2.9k
BFS-V 1 -1.295 -1.295 61.8
BFS-V 10 -1.523 -1.065 0.6k
BFS-V 20 -1.520 -0.948 1.2k
BFS-V 50 -1.474 -0.813 3.1k
DFS-V 1 -1.295 -1.295 61.8
DFS-V 10 -1.498 -1.067 67.8
DFS-V 20 -1.507 -0.985 71.8
DFS-V 50 -1.503 -0.86 85.8
MCTS-$\alpha$ (no sampling) 1 2.221 2.221 186
MCTS-$\alpha$-intra tree 1 1.538 1.538 198.50
MCTS-$\alpha$-intra tree 10 1.527 3.052 1.6k
MCTS-$\alpha$-intra tree 20 1.533 3.311 3.1k
MCTS 1 -1.295 -1.295 61.8
MCTS 10 -1.146 0.160 0.6k
MCTS 20 -1.08 0.528 1.2k
MCTS 50 -0.961 0.981 2.8k
MCTS-Rollout 1 1.925 1.925 0.8k
MCTS-Rollout 10 2.278 2.540 1.1k
MCTS-Rollout 20 2.376 2.643 1.2k
MCTS-Rollout 50 2.491 2.746 1.3k